OpenStudy (anonymous):

(1-x)^-2(x-1) simplify

OpenStudy (anonymous):

(1-x) (1-x) - 2x + 2 = 1 - 2x + x^2 - 2x + 2 = x^2 - 4x + 3 = (x - 3) (x - 1)

OpenStudy (anonymous):

Sorry i meant (1-x)^-2*(x-1)

OpenStudy (anonymous):

[(1/(x-1)^2] * (x-1) = 1/(x^2- 4x + 3) = (x-1) / (x-1) (x-3) = 1/(x-3) =(x-3)^-1